OpenStudy (anonymous):

what is -1 times -2?

OpenStudy (vanillaangelic):

2. Negetive times negetive is positive.

OpenStudy (day00):

2

OpenStudy (anonymous):

Please explain in steps.

OpenStudy (kaylak12345):

negative times negative is positive positive times negative is negative positive times positive is positive

OpenStudy (kaylak12345):

what is 1 times 2

OpenStudy (anonymous):

Thanks! ;)

OpenStudy (anonymous):

3

OpenStudy (kaylak12345):

no you are adding

OpenStudy (kaylak12345):

whats 1 times 1

OpenStudy (anonymous):

Oops! Ahhh! 1 x 1 is 1 and 1 x 2 is 2

OpenStudy (kaylak12345):

yes lol

OpenStudy (kaylak12345):

now anyone here wanna help me in geometry lol

OpenStudy (kaylak12345):

if so ill tag you in my question

OpenStudy (kaylak12345):

and what grade are you in the one

OpenStudy (anonymous):

6th

OpenStudy (kaylak12345):

ahhhh about when I learned this lol

OpenStudy (day00):

ok

OpenStudy (anonymous):

I wish I could help.... but I'm not learning geometry quite yet

OpenStudy (mathstudent55):

This pattern shows that positive times negative is negative. Notice that each product is 1 less than the previous product. What is -1 * 1? 5 * 1 = 5 4 * 1 = 4 3 * 1 = 3 2 * 1 = 2 1 * 1 = 1 0 * 1 = 0 -1 * 1 = ? -2 * 1 = ? -3 * 1 = ?

OpenStudy (kaylak12345):

lol i know the one im taking 10th grade math anyways lol

OpenStudy (mathstudent55):

5 * 1 = 5 4 * 1 = 4 3 * 1 = 3 2 * 1 = 2 1 * 1 = 1 0 * 1 = 0 -1 * 1 = -1 -2 * 1 = -2 -3 * 1 = -3 The last three multiplications follow the pattern above them, and each product is 1 less than the product before.

OpenStudy (anonymous):

Okay thanks! ;)

OpenStudy (mathstudent55):

Now we know negative times positive is negative. What about negative times negative? Here is a new pattern for negative time negative. 5 * (-1) = -5 4 * (-1) = -4 3 * (-1) = -3 2 * (-1) = -2 1 * (-1) = -1 0 * (-1) = 0 -1 * (-1) = ? -2 * (-1) = ? -3 * (-1 ) = ? What are the last three products above? You see that by following the pattern, each product is 1 more than the product before, so the last three products are 1, 2, and 3. That shows that -1 * (-1) = 1, etc., or negative times negative equals positive.

OpenStudy (anonymous):

Wow this really helped thanks! :D

OpenStudy (mathstudent55):

You're welcome.